# Unifying Attention Heads and Task Vectors via Hidden State Geometry in In-Context Learning

### v1

@@ F.1 Replication of subplot (A) of Figure 4 for other datasets and models
In Figure 2025 we provide the visualizations of the experiments presented in subplot (A) of Figure 4, to establish the generalizability of our conclusion that the directional alignment and the encoding of label-related semantics into the hidden states happen almost simultaneously after the phase transition. The layer-wise trends of the directional alignment measures and output alignment exhibit high correlation across all datasets and models.
F.2 Replication of subplot (B) of Figure 4 for other datasets
@@
In Table 3-7, we present the tokens decoded from top singular directions near the phase transition breakpoint for other datasets. The results demonstrate that the sudden .semantic shift toward label-related tokens in the hidden states (as demonstrated by subplot (B) of Figure 4) is present in other datasets as well. Note that the phase transition happens at different layers for different datasets.
F.3 Replication of subplot (B) of Figure 4 with replaced demonstration labels
@@
As shown in Appendix J, the results reveal an interesting pattern. The first major shift in the layer-wise semantic content occurs at layer 24, where the ground-truth label tokens suddenly become dominant. This closely mirrors the pattern observed with normal demonstration labels in Table 5, and aligns with the point where output alignment begins to rise in Figure 20 (C). However, another drastic shift is observed at layer 45, where the decoded top tokens begin to include the symbol tokens. These symbols suddenly appear and start to compete with semantically relevant task tokens, marking the onset of alignment between the hidden states and the unembedding vectors of the symbols. This delayed alignment corresponds with the pattern seen in subplot (C) of Figure 3, where the phase transition under demonstrations with symbolized labels is postponed to later layers.
@@
F.4 Replication of subplot (C) of Figure 4 for other datasets
Subplot (C) of Figure 4 reveals a two-stage pattern in the layer-wise retainment and filtering of semantic information, where most post-transition layers retain label-related semantics and filter out irrelevant ones, and the final layers do the opposite. We show in Table 10-14 that it exists in other datasets besides SST-2 as well.
F.5 Replication of Figure 5 for other datasets and models
We first provide the results of rank-10 approximations on all 6 datasets for Llama2-70B in Figure 39. The conclusion is that low-rank denoising significantly increase the output alignment significantly acorss all datasets, particularly in the innitial layers after the phase transition. This confirms the generality of our findings in Subsection 5.2. Additionally, in the final layers, the output alignment also increases on on SST-2, SUBJ, TREC, and SNLI, though not on RTE and CB.
In Figure 40-45, we present results for other models. These are consistent with the findings from Llama2-70B. Specifically, the pattern that low-rank denoising improves output alignment more in text classification datasets than in natural language inference datasets holds across models. This suggests a higher level of linguistic complexity in natural language inference tasks—likely because such tasks require reasoning over both a premise and a hypothesis, whose combined semantics are more difficult to compress and more prone to information loss during low-rank projection.
F.6 Ablation experiments for the rank hyperparameter in low-rank approximations
In this section, we explore the effect of varying the number of retained ranks in rank-$r$ SVD approximations of Llama2-70B hidden states across all datasets. Specifically, we investigate how increasingly aggressive rank reductions—setting $r=1$, $2$, and $5$—impact performance. The results, shown in Figure 46, Figure 47, and Figure 48, yield several notable findings.
@@ F.6 Ablation experiments for the rank hyperparameter in low-rank approximations
Appendix G Experiment details concerning the identification of PTHs and IHs
@@
Identification of PTHs   For each $x_{i}$, the LLM will generate an attention tensor $\bm{Attn}_{i}\in{\mathbb{R}}^{L\times N_{h}\times s(x_{i})\times s(x_{i})}$, where $L$ is the total number of layers, $N_{h}$ is the number of attention heads per layer. Hence, for the $n_{h}^{th}$ attention head at layer $l$, there is a corresponding attention matrix $\bm{Attn}(l,n_{h})_{i}\in{\mathbb{R}}^{s(x_{i})\times s(x_{i})}$, where $\bm{Attn}(l,n_{h})_{i,j,k}$ is the attention of this attention head at the $k^{th}$ token to the $j^{th}$ token in $x_{i}$. The PTH score of this attention head on $x_{i}$ is thus defined as $\sum_{k=2}^{s(x_{i})-1}\bm{Attn}(l,n_{h})_{i,k-1,k}$, which measures the total attention the attention head assigns at each token position to the previous token. Consequently, the PTH score of the $n_{h}^{th}$ attention head at layer $l$ on all the queries is $\sum_{i=1}^{50}\sum_{k=2}^{s(x_{i})-1}\bm{Attn}(l,n_{h})_{i,k-1,k}$. We calculate the PTH scores for all $(l,n_{h})$ pairs and choose the top 10% attention heads as the identified PTHs.
@@
Appendix H Supplementary materials for Subsection 5.3
H.1 Replication of Figure 6 for other models
In Figure 26-31, we present the results of attention heads ablations across models. The results support our conclusions regarding the significance of PTHs in enhancing the separability of the ICL hidden states and the significance of IHs in aligning them with the label unembedding vectors.
H.2 Results of attention heads ablation in the ICL setting where query label is not in context
The results for other models, reported in Figure 26-31, confirm the generality of our conclusions across model sizes and architectures. To test whether these findings hold under different ICL settings as well, we replicate the attention head ablation experiments in a special ICL scenario (“In-weight Learning”) studied by Reddy [2024] and Chan et al. [2022], where none of the in-context demonstrations share the query’s label. This setting is useful for analyzing attention heads—especially IHs, which are known for copy-like behavior—when direct label copying is not possible.
@@ H.2 Results of attention heads ablation in the ICL setting where query label is not in context
H.3 Results of attention heads ablation in the zero-shot setting
@@
Appendix I Experiment details of using IH outputs to construct task vectors
For each dataset, we first construct 8-shot ICL prompts using the first 50 queries. The demonstrations used are exactly the same as the ones used to evaluate the 8-shot ICL accuracy for each dataset. Then, following the procedures described in Todd et al. [2024], we obtain the average output (across the 50 prompts) of each identified top IH at the final token position, and sum the average outputs of all IHs to obtain the task vector.
@@ Appendix I Experiment details of using IH outputs to construct task vectors
Appendix J Broader impacts
@@
A Varying number of demonstrations
B kNN Demonstration selection
C Random symbols as demonstration labels
@@
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
@@
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
@@
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
@@
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
@@
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
@@
A SUBJ
B TREC
C SNLI
D RTE
E CB
@@ D RTE
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
@@
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
@@
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
@@
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
@@
Layer 79 false, !, ., #, True, <s>, !., Hash, Street, {CJK}UTF8gbsn群
@@
Layer 50 ловi, CA, Um, wob, prot, jär, {CJK}UTF8gbsn望, ang, {CJK}UTF8gbsn孝, Roman
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@ E RTE
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
@@
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB

### v2

@@ G.1 Replication of subplot (A) of Figure 5 for other datasets and models
In Figure 22- 28 we provide the visualizations of the experiments presented in subplot (A) of Figure 5, to establish the generalizability of our conclusion that the directional alignment and the encoding of label-related semantics into the hidden states happen almost simultaneously after the phase transition. The layer-wise trends of the directional alignment measures and output alignment exhibit high correlation across all datasets and models.
G.2 Replication of subplot (B) of Figure 5 for other datasets
@@
In Table 4-8, we present the tokens decoded from top singular directions near the phase transition breakpoint for other datasets. The results demonstrate that the sudden .semantic shift toward label-related tokens in the hidden states (as demonstrated by subplot (B) of Figure 5) is present in other datasets as well. Note that the phase transition happens at different layers for different datasets.
G.3 Replication of subplot (B) of Figure 5 with replaced demonstration labels
@@
As shown in Table 9, the results reveal an interesting pattern. The first major shift in the layer-wise semantic content occurs at layer 24, where the ground-truth label tokens suddenly become dominant. This closely mirrors the pattern observed with normal demonstration labels in Table 6, and aligns with the point where output alignment begins to rise in Figure 22 (C). However, another drastic shift is observed at layer 45, where the decoded top tokens begin to include the symbol tokens. These symbols suddenly appear and start to compete with semantically relevant task tokens, marking the onset of alignment between the hidden states and the unembedding vectors of the symbols. This delayed alignment corresponds with the pattern seen in subplot (C) of Figure 3, where the phase transition under demonstrations with symbolized labels is postponed to later layers.
@@
G.4 Replication of subplot (C) of Figure 5 for other datasets
Subplot (C) of Figure 5 reveals a two-stage pattern in the layer-wise retainment and filtering of semantic information, where most post-transition layers retain label-related semantics and filter out irrelevant ones, and the final layers do the opposite. We show in Table 11-15 that it exists in other datasets besides SST-2 as well.
G.5 Replication of Figure 6 for other datasets and models
We first provide the results of rank-10 approximations on all 6 datasets for Llama2-70B in Figure 42. The conclusion is that low-rank denoising significantly increase the output alignment significantly acorss all datasets, particularly in the innitial layers after the phase transition. This confirms the generality of our findings in Subsection˜5.2. Additionally, in the final layers, the output alignment also increases on on SST-2, SUBJ, TREC, and SNLI, though not on RTE and CB.
In Figure 43-48, we present results for other models. These are consistent with the findings from Llama2-70B. Specifically, the pattern that low-rank denoising improves output alignment more in text classification datasets than in natural language inference datasets holds across models. This suggests a higher level of linguistic complexity in natural language inference tasks—likely because such tasks require reasoning over both a premise and a hypothesis, whose combined semantics are more difficult to compress and more prone to information loss during low-rank projection.
G.6 Ablation experiments for the rank hyperparameter in low-rank approximations
In this section, we explore the effect of varying the number of retained ranks in rank-$r$ SVD approximations of Llama2-70B hidden states across all datasets. Specifically, we investigate how increasingly aggressive rank reductions—setting $r=1$, 22, and 55—impact performance. The results, shown in Figure 49, Figure 50, and Figure 51, yield several notable findings.
@@ G.6 Ablation experiments for the rank hyperparameter in low-rank approximations
Appendix H Experiment details concerning the identification of PTHs and IHs
@@
Identification of PTHs  For each $x_{i}$, the LLM will generate an attention tensor $\bm{Attn}_{i}\in{\mathbb{R}}^{L\times N_{h}\times s(x_{i})\times s(x_{i})}$, where $L$ is the total number of layers, $N_{h}$ is the number of attention heads per layer. Hence, for the $n_{h}^{th}$ attention head at layer $l$, there is a corresponding attention matrix $\bm{Attn}(l,n_{h})_{i}\in{\mathbb{R}}^{s(x_{i})\times s(x_{i})}$, where $\bm{Attn}(l,n_{h})_{i,j,k}$ is the attention of this attention head at the $k^{th}$ token to the $j^{th}$ token in $x_{i}$. The PTH score of this attention head on $x_{i}$ is thus defined as $\sum_{k=2}^{s(x_{i})}\bm{Attn}(l,n_{h})_{i,k-1,k}$, which measures the total attention the attention head assigns at each token position to the previous token. Consequently, the PTH score of the $n_{h}^{th}$ attention head at layer $l$ on all the queries is $\sum_{i=1}^{50}\sum_{k=2}^{s(x_{i})}\bm{Attn}(l,n_{h})_{i,k-1,k}$. We calculate the PTH scores for all $(l,n_{h})$ pairs and choose the top 10% attention heads as the identified PTHs.
@@
Appendix I Supplementary materials for Subsection˜5.3
I.1 Replication of Figure 7 for other models
In Figure 29-34, we present the results of attention heads ablations across models. The results support our conclusions regarding the significance of PTHs in enhancing the separability of the ICL hidden states and the significance of IHs in aligning them with the label unembedding vectors.
I.2 Replication of Figure 8 for other models
In Figure 59-64, we present the results of analyzing the geometric contribution of PTHs and IHs in other models to the evolution of hidden states by injecting their outputs as task vectors. The results clearly demonstrate their respective significance in fostering the separability and alignment of hidden states. In particular, the PTH task vectors of other models exhibit greater potential to increase hidden-state separability than the result shown in Figure 8.
I.3 Replication of Table 1 for other models
In Table 16-21, we compare the layer distributions of the top 1%, 2%, 5%, and 10% identified IHs and PTHs in other models and conduct Mann–Whitney U tests to assess the significance of the differences. The results demonstrate a consistent difference in the layer positions of PTHs and IHs across models, with PTHs significantly preceding IHs on average—the difference being most prominent in the two Gemma-family models, Gemma-2B and Gemma-7B.
I.4 Replication of Figure 9 for other models
In Figure 67-70, we show the results of injecting task vectors derived from IH outputs for other models. The task vectors are similarly injected at layer $\frac{3}{8}L$, where $L$ is the total number of layers of the respective models. The results demonstrate that IH outputs qualify as effective task vectors across models, with the exception of Gemma-2B, as they considerably improve prediction accuracy in the zero-shot setting after being injected into the zero-shot hidden states.
I.5 Results of attention heads ablation in the ICL setting where query label is not in context
The results for other models, reported in Figure 29-34, confirm the generality of our conclusions across model sizes and architectures. To test whether these findings hold under different ICL settings as well, we replicate the attention head ablation experiments in a special ICL scenario (“In-weight Learning”) studied by Reddy [2024] and Chan et al. [2022], where none of the in-context demonstrations share the query’s label. This setting is useful for analyzing attention heads—especially IHs, which are known for copy-like behavior—when direct label copying is not possible.
@@ I.5 Results of attention heads ablation in the ICL setting where query label is not in context
I.6 Results of attention heads ablation in the zero-shot setting
@@
Appendix J Experiment details of using IH outputs to construct task vectors
For each dataset, we first construct 8-shot ICL prompts using the first 50 queries. The demonstrations used are exactly the same as the ones used to evaluate the 8-shot ICL accuracy for each dataset. Then, following the procedures described in Todd et al. [2024], we obtain the average output (across the 50 prompts) of each identified top 10% IHs and PTHs at the final token positions, and sum the average outputs of all IHs to obtain the task vector.
@@
Appendix K Broader impacts
@@
(a) Varying number of demonstrations
(b) kNN Demonstration selection
(c) Random symbols as demonstration labels
@@
(a) Varying number of demonstrations
(b) KNN Demonstration selection
(c) Random symbols as demonstration labels
@@
(a) Varying number of demonstrations
(b) KNN Demonstration selection
(c) Random symbols as demonstration labels
@@
(a) Varying number of demonstrations
(b) KNN Demonstration selection
(c) Random symbols as demonstration labels
@@
(a) Varying number of demonstrations
(b) KNN Demonstration selection
(c) Random symbols as demonstration labels
@@
(a) Varying number of demonstrations
(b) KNN Demonstration selection
(c) Random symbols as demonstration labels
@@
(a) SUBJ
(b) TREC
(c) SNLI
(d) RTE
(e) CB
@@
(a) SUBJ
(b) TREC
(c) SNLI
(d) RTE
(e) CB
(f) SST-2
(a) SUBJ
(b) TREC
(c) SNLI
(d) RTE
(e) CB
(f) SST-2
(a) SUBJ
(b) TREC
(c) SNLI
(d) RTE
(e) CB
(f) SST-2
@@
(a) SUBJ
(b) TREC
(c) SNLI
(d) RTE
(e) CB
(f) SST-2
@@
(a) SUBJ
(b) TREC
(c) SNLI
(d) RTE
(e) CB
(f) SST-2
@@
(a) SUBJ
(b) TREC
(c) SNLI
(d) RTE
(e) CB
(f) SST-2
Figure 28: Output and directional alignment of Gemma-7B ICL hidden states with label unembedding vectors on various datasets.
@@
Layer 79 false, !, ., #, True, <s>, !., Hash, Street, 群
@@
Layer 50 ловi, CA, Um, wob, prot, jär, 望, ang, 孝, Roman
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
(a) SST-2
(b) SUBJ
(c) TREC
(d) SNLI
(e) RTE
(f) CB
@@
Table 16: Mean layer values of IH and PTH along with p-values of Mann-Whitney U test on Llama3-8B.
% Level IH mean layer PTH mean layer p-value of Mann-Whitney U test
1% 15.883 9.5667 $5.06\times 10^{-7}$
2% 16.392 12.658 $3.99\times 10^{-5}$
5% 16.493 13.833 $2.29\times 10^{-5}$
10% 16.435 14.364 $4.46\times 10^{-6}$
Table 17: Mean layer values of IH and PTH along with p-values of Mann-Whitney U test on Llama3-70B.
% Level IH mean layer PTH mean layer p-value of Mann-Whitney U test
1% 47.624 45.755 0.5573
2% 46.273 46.977 0.1347
5% 44.354 47.277 $1.18\times 10^{-7}$
10% 43.278 44.257 $8.38\times 10^{-4}$
Table 18: Mean layer values of IH and PTH along with p-values of Mann-Whitney U test on Llama2-7B.
% Level IH mean layer PTH mean layer p-value of Mann-Whitney U test
1% 15.883 8.5833 $5.62\times 10^{-7}$
2% 16.133 10.108 $9.78\times 10^{-9}$
5% 16.301 12.092 $4.12\times 10^{-8}$
10% 16.405 12.559 $7.27\times 10^{-13}$
Table 19: Mean layer values of IH and PTH along with p-values of Mann-Whitney U test on Llama2-13B.
% Level IH mean layer PTH mean layer p-value of Mann-Whitney U test
1% 21.312 17.094 $1.75\times 10^{-2}$
2% 20.135 18.026 $9.44\times 10^{-2}$
5% 19.960 16.927 $6.08\times 10^{-6}$
10% 19.888 16.343 $3.08\times 10^{-13}$
Table 20: Mean layer values of IH and PTH along with p-values of Mann-Whitney U test on Gemma-2B.
% Level IH mean layer PTH mean layer p-value of Mann-Whitney U test
1% 9.0000 1.6667 $8.58\times 10^{-3}$
2% 10.917 5.0000 $7.91\times 10^{-4}$
5% 11.571 4.4286 $2.22\times 10^{-10}$
10% 12.214 6.1667 $2.48\times 10^{-12}$
Table 21: Mean layer values of IH and PTH along with p-values of Mann-Whitney U test on Gemma-7B.
% Level IH mean layer PTH mean layer p-value of Mann-Whitney U test
1% 18.708 9.0833 $2.01\times 10^{-5}$
2% 19.062 9.1250 $1.96\times 10^{-6}$
5% 18.742 7.6212 $6.31\times 10^{-21}$
10% 18.523 9.0530 $4.44\times 10^{-32}$
(a) Layer-wise separability score: 0-shot vs. 8-shot
(b) Layer-wise composite alignment: 0-shot vs. 8-shot
Figure 53: Layer-wise trends of separability score and composite alignment of 0-shot and 8-shot hidden states of Llama3-8B.
(a) Layer-wise separability score: 0-shot vs. 8-shot
(b) Layer-wise composite alignment: 0-shot vs. 8-shot
Figure 54: Layer-wise trends of separability score and composite alignment of 0-shot and 8-shot hidden states of Llama3-70B.
(a) Layer-wise separability score: 0-shot vs. 8-shot
(b) Layer-wise composite alignment: 0-shot vs. 8-shot
Figure 55: Layer-wise trends of separability score and composite alignment of 0-shot and 8-shot hidden states of Llama2-7B.
(a) Layer-wise separability score: 0-shot vs. 8-shot
(b) Layer-wise composite alignment: 0-shot vs. 8-shot
Figure 56: Layer-wise trends of separability score and composite alignment of 0-shot and 8-shot hidden states of Llama2-13B.
(a) Layer-wise separability score: 0-shot vs. 8-shot
(b) Layer-wise composite alignment: 0-shot vs. 8-shot
Figure 57: Layer-wise trends of separability score and composite alignment of 0-shot and 8-shot hidden states of Gemma-2B.
(a) Layer-wise separability score: 0-shot vs. 8-shot
(b) Layer-wise composite alignment: 0-shot vs. 8-shot
Figure 58: Layer-wise trends of separability score and composite alignment of 0-shot and 8-shot hidden states of Gemma-7B.
(a) Layer-wise separability score: PTH task vector
(b) Layer-wise composite alignment: IH task vector
Figure 59: Effects of injecting PTH and IH outputs as task vectors on the geometric properties of Llama3-8B zero-shot hidden states.
(a) Layer-wise separability score: PTH task vector
(b) Layer-wise composite alignment: IH task vector
Figure 60: Effects of injecting PTH and IH outputs as task vectors on the geometric properties of Llama3-70B zero-shot hidden states.
(a) Layer-wise separability score: PTH task vector
(b) Layer-wise composite alignment: IH task vector
Figure 61: Effects of injecting PTH and IH outputs as task vectors on the geometric properties of Llama2-7B zero-shot hidden states.
(a) Layer-wise separability score: PTH task vector
(b) Layer-wise composite alignment: IH task vector
Figure 62: Effects of injecting PTH and IH outputs as task vectors on the geometric properties of Llama2-13B zero-shot hidden states.
(a) Layer-wise separability score: PTH task vector
(b) Layer-wise composite alignment: IH task vector
Figure 63: Effects of injecting PTH and IH outputs as task vectors on the geometric properties of Gemma-2B zero-shot hidden states.
(a) Layer-wise separability score: PTH task vector
(b) Layer-wise composite alignment: IH task vector
Figure 64: Effects of injecting PTH and IH outputs as task vectors on the geometric properties of Gemma-7B zero-shot hidden states.
Figure 65: Average effect across datasets of steering Llama3-8B zero-shot hidden states using task vectors created from IH outputs.
Figure 66: Average effect across datasets of steering Llama3-70B zero-shot hidden states using task vectors created from IH outputs.
Figure 67: Average effect across datasets of steering Llama2-7B zero-shot hidden states using task vectors created from IH outputs.
Figure 68: Average effect across datasets of steering Llama2-13B zero-shot hidden states using task vectors created from IH outputs.
Figure 69: Average effect across datasets of steering Gemma-2B zero-shot hidden states using task vectors created from IH outputs.
Figure 70: Average effect across datasets of steering Gemma-7B zero-shot hidden states using task vectors created from IH outputs.
Table 22: Comparison of classification accuracy: training classifier v.s. direct decoding with zero-shot prompts.
Model 0-shot (%) 0-shot+classifier (%) Improvement
Llama3-8B 0.30% 73.71% +73.41%
Llama3-70B 0.00% 78.97% +78.97%
Llama2-7B 0.31% 75.63% +75.32%
Llama2-13B 0.36% 79.71% +79.35%
Llama2-70B 0.24% 79.02% +78.78%
Gemma-2B 0.00% 75.56% +75.56%
Gemma-7B 0.02% 75.43% +75.41%
Table 23: Comparison of classification accuracy: training classifier v.s. direct decoding with zero-shot prompts.
Model 8-shot (%) 8-shot+classifier (%) Improvement
Llama3-8B 76.20% 85.47% +9.27%
Llama3-70B 81.01% 84.46% +3.45%
Llama2-7B 74.83% 85.59% +10.76%
Llama2-13B 74.82% 85.44% +10.62%
Llama2-70B 80.52% 87.76% +7.24%
Gemma-2B 67.03% 77.87% +10.84%
Gemma-7B 79.19% 84.94% +5.75%
Table 24: Comparison of classification accuracy: regular ICL v.s. low-rank denoising of hidden states
Model 8-shot (%) 8-shot+denoising (%) Improvement
Llama3-8B 76.20% 78.58% +2.38%
Llama3-70B 81.01% 82.40% +1.39%
Llama2-7B 74.83% 77.68% +2.85%
Llama2-13B 74.82% 79.52% +4.70%
Llama2-70B 80.52% 80.72% +0.20%
Gemma-2B 67.03% 67.25% +0.22%
Gemma-7B 79.19% 66.11% -13.08%